A couch is 8 1/6 long. The material needed to make a slipcover is to be 2 1/3 times as long to allow for hems and tucking under. How long a piece of material needed?
Answer by checkley71(8403)

8 1/6*2 1/3=49/6*7/3=343/18=19 1/18 ANSWER.